In stationary wave the overtone shortcut formula is. ..,...? Plz give me that

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
Zeroth overtone is fundamental F=v/4L , lamda=4L
First overtone is 3th harmonic F2=3F, lambda2 =2/3*lamda
Second is 5th harmonic …..

Continuing like this n-1/2 is F_(n-1)/2=nF, lambda_(n-1)/2=2lambda/n
Thanks & Regards
Arun Kumar
Btech IIT delhi
Askiitians Faculty